OpenStudy (anonymous):

Hello, I had a test yesterday and I need your help to understand some questions. 1. is it possible that a matrix A^3 = I , but A != I (over M2x2(R)). The answer is yes but I can't seem to find such a matrix. Notice your matrix value must be real values and cannot be complex numbers or something else. 2. If A and B both matrix with the same order nxn, and X is also nxn matrix and invertible, and AX=B then rank(A)=rank(B). I guess that because X is invertible its rank doesn't change the rank of A. Am I right? 3. If A is a nxn matrix then det(A - A^t) = 0.

OpenStudy (anonymous):

Regarding the 3rd question : I'm not sure if its correct. any example of matrix I'm trying to use does satisfies this equation. Could someone tell me if it's ture and if it is true tell me how to prove it please. Thank you!!!!

OpenStudy (unklerhaukus):

A != I what does this mean

OpenStudy (anonymous):

i guess A is not the identity matrix

OpenStudy (anonymous):

It means that your A matrix isn't the Identity matrix. I = Identity matrix

OpenStudy (unklerhaukus):

so != means is not equal to

OpenStudy (anonymous):

yes @Coolsector is right.

OpenStudy (anonymous):

yep. A isn't the identity matrix. what is it then...

OpenStudy (unklerhaukus):

i found this solution \[\mathbf A=\left(\begin{array}{cc} (-1)^{2/3}&0\\ 0&(-1)^{2/3}\end{array}\right)\] but \((-1)^{2/3}\) is complex

OpenStudy (anonymous):

yeah. you can't use complex. that's why it's so wierd.

OpenStudy (anonymous):

3. If A is a nxn matrix then det(A - A^t) = 0. this is not correct

OpenStudy (anonymous):

as for the rank.. there is a property Rank(AB) <= Min(Rank(A),Rank(B)) since x is invertible its rank is so Rank(Ax)<=Rank(A) and it means RanK(B) <= Rank(A) but we can multiply both sides of Ax=B by x-1 and get A = Bx-1 so Rank(Bx-1) <= Rank(B) since rank(x-1) = n and it means Rank(A) <= Rank(B) so we have Ran(B) <= Rank(A) Rank(A) <= Rank(B) -> Rank(A) = Rank(B)

OpenStudy (anonymous):

WHY

OpenStudy (anonymous):

First of all thank you very much #Coolsector

OpenStudy (anonymous):

But Why If A is a nxn matrix then det(A - A^t) = 0. is not correct?

OpenStudy (anonymous):

I forgot to mention the n is odd

OpenStudy (anonymous):

I don't know if it changes something.

OpenStudy (anonymous):

oh so let me check since i checked one even n and get it wrong .. as for the rank you understand my answer ?

OpenStudy (anonymous):

I'm trying to understand it, just a second. I'm sorry I forogt to mention about the n odd....

OpenStudy (anonymous):

Yeah I think I understand the rank proof. thank you!

OpenStudy (anonymous):

for the odd n it looks correct .. i see the pattern but not sure how to come with a proof

OpenStudy (anonymous):

Okay, thank you very much for you help! I'll think about it more. There must be a way to prove it beccause it always work

OpenStudy (anonymous):

yes im sure.. i just dont know how :(

OpenStudy (anonymous):

you see you always left with a matrix with zero in the diagonal and it is an anti symmetric matrix

OpenStudy (anonymous):

so if we have anti symmetric matrix A-A^T = - (A-A^T)^T

OpenStudy (anonymous):

now lets call A-A^T = B so from what i wrote we get B = -B^T so Det(B) = Det(-B^T) = (-1)^n * Det(B^T) but for every matrix Det(B) = Det(B^T) now for an odd n we have Det(B) = -Det(B) -> Det(B) = 0

OpenStudy (anonymous):

now the only thing you have to do is to show that A-A^T is always an anti symmetric matrix

OpenStudy (anonymous):

but im happy with this proof

OpenStudy (anonymous):

@mtaOS

OpenStudy (anonymous):

You are good man!! thank you!